# LightCode: Light Analytical and Neural Codes for Channels with Feedback

### v2

@@ VI-B Training
Our primary region of interest is rate $\nicefrac{{3}}{{9}}$ code at SNR $-1.0$ dB, where our target BLER is $\approx 10^{-9}$. This is the current state-of-the-art performance by GBAF [26]. In order to achieve such an extremely low error rate, it is important to reliably train and simulate large amounts of samples. Inspired by the high-SNR variant of Gallager-Nakibŏglu [4], we hypothesize that in regions of high-SNR or low errors, a significant benefit to the forward transmission arises from the power reallocation to symbols with non-zero errors. This can only be realized during the training by considering a large batch and enforcing the power constraint per batch. Moreover, it is well understood that deep-learning models generalize better when trained with large batch sizes. Hence, we consider an extremely large batch size of $10^{5}$. While this is significantly larger compared to GBAF, which has a batch size of 8192, it is still a smaller number of symbols, considering that GBAF contains 17 symbols per codeblock, whereas our scheme consists of only 1 symbol per codeblock. Further, it is possible for us to train with such a large batch size because of the small number of parameters compared to other deep-learning-based coding schemes. For a fair comparison, we follow a training methodology similar to that of GBAF, as explained below.
Input: Encoder model $\phi$, Decoder model $\psi$, Block length $K$, number of rounds $D$, noise variance $\sigma_{ff}^{2}$, batch size $B$, number of epochs $E$, learning rate lr
for $i\leq E$ do
       Generate batch of random binary vectors $\mathbf{u}\in\{0,1\}^{K\times B}$
       for $i\leq D$ do
             $x_{i}=\phi\left(\mathbf{u},y_{1},y_{2},\dots,y_{i-1}\right)$ ;
@@
             $y_{i}=x_{i}+n_{i}$
       end for
@@
       Compute the NLL loss $\frac{1}{B}\sum_{j=1}^{B}L_{\text{NLL}}(\mathbf{c_{j}},\mathbf{p_{u_{j}}})$, where $\mathbf{c_{j}}$ is the class index corresponding to $j^{\text{th}}$ message vector $\mathbf{u_{j}}$ and $\mathbf{p_{u_{j}}}$ is class probability vector after the SoftMax layer.
       Clip the gradients to $0.5$.
       Update model parameters for $\phi$ and $\psi$ using AdamW optimizer with learning rate lr.
       Update the learning rate using LambdaLR.
end for
@@
We use AdamW optimizer, a stochastic optimization method that modifies the typical implementation of weight decay in Adam by decoupling weight decay from the gradient update. We initialize the learning to $10^{-3}$ and use a LambdaLR scheduler with a weight decay of $0.01$. Additionally, we clip all the gradient values to $0.5$ for numerical stability. Starting with randomly initialized weights, we train the encoder and decoder models simultaneously on $1.2\times 10^{5}$ batches. Each input symbol corresponds to $K$ bits, resulting in a $2^{K}$ category classification problem for the decoder. Accordingly, we use cross entropy (CE) loss to measure the performance as
$L_{\text{CE}}=\frac{1}{B}\sum_{i=1}^{B}\left(\sum_{j=0}^{2^{K}-1}c_{ij}\log p_{ij}\right),$
@@
Optimizer AdamW
TABLE I: Hyperparameters used in model training for rate $\nicefrac{{3}}{{9}}$, SNR -1.0 LightCode.
@@ VII Main Results
We begin by comparing the performance of LightCode with the current state-of-the-art in deep-learning-based coding schemes for noiseless passive feedback setting, GBAF [26] and other schemes, demonstrated in Fig. 5. Next, we discuss a method for extending LightCode to moderate block-length regimes of up to $L=300$ and study the performance.
@@
In this section, we evaluate the performance of LightCode and PowerBlast and compare them against several existing analytical and deep-learning feedback schemes. For concreteness, we consider the canonical setting of rate $R=\nicefrac{{3}}{{9}}$ with block length $K=3$ and a number of rounds $D=9$ transmitted on AWGN forward channel and noiseless feedback.
Baselines. Our primary comparison is against GBAF [26], which is the current state-of-the-art for the noiseless passive feedback setting. Additionally, we consider Deepcode [21], DEFC [22], DRFC [23], Attentioncode [25], and Robust Coding [21]. Additionally, for completeness as well as to understand the relative gains of deep-learning-based coding schemes, we also compare the performance against Schalkwijk-Kailath [2], Gallager-Nakibŏglu [4] and PowerBlast.
In Fig. 5, we compare the BLER performance of rate $\nicefrac{{3}}{{9}}$ coding schemes. LightCode consistently outperforms the existing deep-learning-based schemes, including GBAF, while utilizing $<\nicefrac{{1}}{{10}}^{\text{th}}$ the number of parameters. Moreover, these results indicate that PowerBlast surpasses the performance of all existing schemes, including LightCode, when an adequately high signal-to-noise ratio is provided, which is $-0.5$ dB for rate $\nicefrac{{3}}{{9}}$. However, at lower SNRs, the performance of deep-learning codes is significantly better.
Figure 5: Performance comparison of LightCode with GBAF and other existing neural feedback codes for rate $\nicefrac{{3}}{{9}}$ and noiseless feedback. LightCode achieves superior BLER performance compared to GBAF while utilizing $<\nicefrac{{1}}{{10}}^{\text{th}}$ the number of parameters.
Multiple rates. The coding rate of LightCode is determined by two factors, block length $K$ and number of rounds of communication $D$. We now keep the block length constant and vary $D$ to compare the performance against GBAF across multiple rates, as shown in Table II. LightCode consistently outperforms GBAF across a range of rates $\{\nicefrac{{3}}{{8}},\nicefrac{{3}}{{7}},\nicefrac{{3}}{{6}},\nicefrac{{3}}{{5}}\}$, while utilizing a fraction of the number of parameters.
@@
TABLE II: BLER performance comparison of LightCode with GBAF with PowerBlast for different rates. LightCode consistently performs better than GBAF while using $<\nicefrac{{1}}{{10}}^{\text{th}}$ the number of parameters.
@@ VII-B Performance in moderate blocklength regime
Using a code designed for $K=3$, it is straightforward to encode larger blocks of length $L$ by treating the message as a packet of $l=\nicefrac{{L}}{{K}}$ independent symbols. Hence, the packet error rate (PER) for a packet of size $l$ can be considered BLER for blocklength $L$. This modular approach to encoding is motivated by two key observations. The first is from the ablation studies in Section VIII-A, which demonstrate that explicit block coding is not beneficial in this regime and does not scale well. Second is the increase in computational complexity of both encoder and decoder modules associated with block coding. Results from Fig. 6 illustrate that LightCode scales gracefully with blocklength by independently encoding information in sets of $K=3$, keeping the complexity constant. Further, we also note that the bit error rate (BER) remains constant regardless of the blocklength $L$.
Figure 6: deep-learning-based coding schemes scale gracefully with the length, providing a flexible way to code for any length without an increase in computational complexity.
@@ VII-C Coding for Channels with Noisy Feedback
Finally, we now consider the case of channels with noisy feedback i.e., $\sigma_{fb}^{2}>0$. While the assumption of noiseless feedback is easier to study, most practical feedback channels suffer from noise even when the receiver sending the feedback operates at high power. Because of this limitation, both SK and GN schemes do not perform well. To address this, [6] introduces a linear feedback scheme that is implemented as an inner code to a concatenated code, which was found to be optimal within the linear family of codes [28]. More recently, [7] proposed using dynamic programming to improve the performance, which turns out to be a generalized version of SK. However, despite these improvements, the linearity of these schemes severely limits the performance that can be achieved.
Recalling the architecture of deep-learning-based schemes introduced in Section V, one of the interesting properties of these codes is their robustness to noise in the feedback channel. By taking advantage of the general architecture, it is straightforward to inject noise into the feedback channel during training to make the encoder-decoder robust. Hence, learning-based schemes can help design a practically realizable class of codes that can trained purely in a data-driven fashion and can provide gains under both noiseless and noisy feedback settings.
Figure 7: Performance comparison of LightCode with GBAF and other existing neural feedback codes for rate $\nicefrac{{3}}{{9}}$ and noisy feedback with SNR = $20$ dB. LightCode achieves BLER performance comparable to GBAF while utilizing only $\nicefrac{{1}}{{10}}^{\text{th}}$ the number of parameters.
@@ VIII Analysis
As evident from results in Section VII, symbol-by-symbol neural schemes can achieve performance comparable to block coding schemes such as GBAF. Consequently, we examine GBAF and LightCode in greater detail to systematically analyze their architecture, training process, and performance, with the aim of identifying the key factors contributing to their performance. Furthermore, we compare the memory and computational complexity of LightCode with existing schemes to quantify the gains. Finally, we provide an interpretation of the encoded representations of LightCode.
@@ VIII-A Ablation studies on GBAF
We investigate the contribution of the self-attention mechanism to the performance of GBAF by performing an ablation study to understand the compute vs performance trade-off better. First, the self-attention mechanism is disabled, and it is observed that this has no significant effect on the BLER performance. Next, both self-attention and positional encoding blocks are disabled, and the performance remains approximately the same. These results, provided in Table III, demonstrate a surprising observation that the self-attention and PE modules, which are responsible for cross-symbol block coding, contribute only marginally to the performance of GBAF. This brings into question the value of block encoding and motivates us to find simpler designs that perform only symbol-by-symbol coding while providing competing performance.
@@
TABLE III: Effect of block coding on performance of GBAF for rate $\nicefrac{{3}}{{9}}$ code with noiseless feedback. Positional encoding and self-attention modules have no noticeable effect on BLER performance.
@@ VIII-B Scaling laws: batch size
In section Section VII, we have demonstrated that it is possible to achieve performance comparable to GBAF with fewer parameters and lower compute complexity. We also hypothesized in Section VI-B that it is important to train the encoder using a very large batch size in order to accurately capture the statistics of the distribution so that the available power can be optimally allocated to the symbols with error at the receiver while reducing the power to the remaining symbols in a batch. In deep learning literature, it is well known that increasing the batch size can noticeably improve the performance of the neural network model [29]. To better understand the effect of batch size on LightCode, we perform a systematic study by training LightCode with different batch sizes. As noted in Fig. 8, at a batch size of $1.5\times 10^{3}$, LightCode has a BLER of $3.7\times 10^{-9}$, similar to that of GBAF (w/o BU). However, the performance drops to $4\times 10^{-10}$, exceeding GBAF, when the batch size increases to $5\times 10^{4}$ and beyond.
@@
Using a lightweight network with a small number of parameters makes it feasible to train with very large batch size, resulting in a significantly lower error rate.
@@ VIII-C Complexity and Throughput
Scheme # parameters
@@
TABLE IV: Rate $\nicefrac{{3}}{{9}}$ LightCode requires $<\nicefrac{{1}}{{10}}^{\text{th}}$ the number of parameters compared to GBAF and Robust Coding.
We compare the total number of parameters in the encoder and decoder for a rate $\nicefrac{{3}}{{9}}$ LightCode against GBAF and Robust Coding within Table IV. LightCode reduces the parameter count by more than 10×10$\times$10 × compared to GBAF and Robust Coding, resulting in significant savings in memory.
@@
TABLE V: Throughput (symbols/sec) comparison. Rate $\nicefrac{{3}}{{9}}$ LightCode achieves up to 10×10$\times$10 × higher decoding throughput compared to GBAF and up to 171×171$\times$171 × higher decoding throughput in CPU mode compared to Robust Coding.
To compare the throughputs, we measure the inference time in CPU mode on a AMD Ryzen Threadripper PRO 5975WX 32-Cores processor as well as in GPU mode using an NVIDIA GeForce RTX 4090, using a batch size of $10^{5}-10^{7}$. LightCode provides up to 10×10$\times$10 × higher throughput compared to GBAF. Further, LightCode provides up to 171×171$\times$171 × higher decoding throughput in CPU mode and up to 10×10$\times$10 × higher throughput in GPU mode compared to Robust Coding, providing significant gains in latency, as shown in Table V.

### v3

@@ VI-B Training
Our primary region of interest is rate $\nicefrac{{3}}{{9}}$ code at SNR $-1.0$ dB, where our target BLER is $\approx 10^{-9}$. This is the current state-of-the-art performance by GBAF [26]. To achieve such an extremely low error rate, it is important to train and simulate large amounts of samples reliably. Inspired by the high-SNR variant of Gallager-Nakiboğlu [4], we hypothesize that in regions of high-SNR or low errors, a significant benefit to the forward transmission arises from the power reallocation to symbols with non-zero errors. This can only be realized during the training by considering a large batch and enforcing the power constraint per batch. Moreover, it is well understood that deep-learning models generalize better when trained with large batch sizes. Hence, we consider an extremely large batch size of $10^{5}$. While this is significantly larger than GBAF, which has a batch size of 8192, it is still a smaller number of symbols, considering that GBAF contains 17 symbols per codeblock. In contrast, our scheme consists of only $1$ symbol per codeblock. Further, it is possible to train with such a large batch size because of the small number of parameters compared to other deep-learning-based coding schemes. For a fair comparison, we follow a training methodology similar to that of GBAF, as explained below.
1
Input: Encoder model $\phi$, Decoder model $\psi$, Block length $K$, number of rounds $D$, forward noise variance $\sigma_{ff}^{2}$, feedback noise variance $\sigma_{fb}^{2}$, batch size $B$, number of epochs $E$, learning rate lr
2 for $i\leq E$ do
3       Generate batch of random binary vectors $\mathbf{u}\in\{0,1\}^{K\times B}$
4       for $i\leq D$ do
@@ 3       Generate batch of random binary vectors $\mathbf{u}\in\{0,1\}^{K\times B}$
5
6            if $\sigma_{fb}^{2}==0$ then
7                   $x_{i}=\phi\left(\mathbf{u},y_{1},y_{2},\dots,y_{i-1},0,\dots,0\right)$
8
9            else
10                   $x_{i}=\phi\left(\mathbf{u},x_{1},\dots,x_{i},n_{1}+\tilde{n}_{1},\dots n_{i}+\tilde{n}_{i-1},\dots,0\right)$
11            $y_{i}=x_{i}+n_{i}$
12
@@ 11            $y_{i}=x_{i}+n_{i}$
13       Compute the multi-class cross entropy loss $\frac{1}{B}\sum_{j=1}^{B}\mathcal{L}_{\text{CE}}(\mathbf{c_{j}},\mathbf{p_{u_{j}}})$, $\mathbf{c_{j}}$ is the class index corresponding to $j^{\text{th}}$ message vector $\mathbf{u_{j}}$ and $\mathbf{p_{u_{j}}}$ is class probability vector after the SoftMax layer.
14       Clip the gradients to $0.5$.
15       Update model parameters for $\phi$ and $\psi$ using AdamW optimizer with learning rate lr.
16       Update the learning rate using LambdaLR.
17
@@
We use AdamW optimizer, a stochastic optimization method that modifies the typical implementation of weight decay in Adam by decoupling weight decay from the gradient update. We initialize the learning to $10^{-3}$ and use a LambdaLR scheduler with a weight decay of $0.01$. Additionally, we clip all the gradient values to $0.5$ for numerical stability. Starting with randomly initialized weights, we jointly train the encoder and decoder models on $1.2\times 10^{5}$ batches. Each input symbol corresponds to $K$ bits, resulting in a $2^{K}$ category classification problem for the decoder. Accordingly, we use the multi-class cross entropy (CE) loss to measure the performance as
$\mathcal{L}_{\text{CE}}=\frac{1}{B}\sum_{i=1}^{B}\left(\sum_{j=0}^{2^{K}-1}c_{ij}\log p_{ij}\right),$
@@
Learning rate scheduler LambdaLR
TABLE I: Hyperparameters for training rate $\nicefrac{{3}}{{9}}$ LightCode.
@@ VII Main Results
We begin by comparing the performance of LightCode with the current state-of-the-art in deep-learning-based coding schemes for noiseless passive feedback setting, GBAF [26] and other schemes, demonstrated in Fig. 5. Next, we discuss a method for extending LightCode to moderate block-length regimes of up to $L=51$ and study the performance.
@@
In this section, we evaluate the performance of LightCode and PowerBlast and compare them against several existing analytical and deep-learning feedback schemes. For concreteness, we consider the canonical setting of rate $R=\nicefrac{{3}}{{9}}$ with block length $K=3$ and $D=9$ rounds of communication on AWGN forward channel and noiseless feedback.
Baselines. Our primary comparison is against GBAF [26], which is the current state-of-the-art for the noiseless passive feedback setting. Additionally, we consider Deepcode [21], DEFC [22], DRFC [23], Attentioncode [25], and Robust Coding [21]. Further, for completeness as well as to understand the relative gains of deep-learning-based coding schemes, we also compare the performance against NR-LDPC [28], Schalkwijk-Kailath [2], Gallager-Nakiboğlu [4] and PowerBlast.
Results. In Fig. 5, we compare the BLER performance of rate $\nicefrac{{3}}{{9}}$ coding schemes. We train a pair of encoder-decoder models at each SNR point in the plot. LightCode consistently outperforms the existing deep-learning-based schemes, including GBAF, while utilizing $<\nicefrac{{1}}{{10}}^{\text{th}}$ the number of parameters. Interestingly, these results indicate that PowerBlast surpasses the performance of all existing schemes, including LightCode, when an adequately high signal-to-noise ratio is provided, which is $-0.5$ dB for rate $\nicefrac{{3}}{{9}}$. However, the performance of deep-learning codes is significantly better at lower SNRs. Here, we highlight that by utilizing the clean feedback, LightCode and PowerBlast achieve an extremely small error rate even when operating at $0.75$ dB below the channel capacity. Further, we note that because of the extremely short block lengths considered, it is hard to derive meaningful achievability or converse bounds, such as [29], where blocks lengths of $100$ or higher are considered.
The blocklengths considered for the baselines in Fig. 5 are listed in Table II, where $K=50$ for some of the schemes, which is larger than the length $K=3$ considered for rest of the schemes including LightCode. While the BLER comparison against the different blocklengths is unfair, to maintain consistency, we used the same methodology followed in the current state-of-the-art results, including GBAF [26] and Robust Coding [24]. For a fairer comparison, we propose a modular approach for extending LightCode to larger blocklengths in Section VII-B.
Scheme Message length $K$
Deepcode $50$
DEFC $50$
DRFC $50$
AttentionCode $50$
Gallager-Nakiboğlu $3$
PowerBlast $3$
Robust Coding $3$
GBAF 33331112
LightCode $3$
TABLE II: Block lengths of different coding schemes.
While GBAF uses a blocklength of $51$, the BLER in [26] was reported for the sub-block of length $3$.
Figure 5: Noiseless feedback: Performance comparison against existing classical and neural feedback codes for rate $\nicefrac{{3}}{{9}}$. PowerBlast achieves the best performance among existing classical schemes and performs comparable to state-of-the-art neural coding schemes in high-SNR regions. LightCode achieves superior BLER performance compared to GBAF while utilizing $<\nicefrac{{1}}{{10}}^{\text{th}}$ the number of parameters.
Multiple rates. The coding rate of LightCode is determined by two factors, block length $K$ and number of rounds of communication $D$. We keep the block length constant and vary $D$ to compare the performance against GBAF across multiple rates, as shown in Table III. LightCode consistently outperforms GBAF across a range of rates $\{\nicefrac{{3}}{{9}},\nicefrac{{3}}{{8}},\nicefrac{{3}}{{7}},\nicefrac{{3}}{{6}},\nicefrac{{3}}{{5}}\}$, while utilizing a fraction of the number of parameters.
@@
TABLE III: BLER performance comparison of LightCode with GBAF with PowerBlast for different rates. LightCode consistently performs better than GBAF while using $<\nicefrac{{1}}{{10}}^{\text{th}}$ the number of parameters.
Error floor. As observed in Fig. 5, LightCode exhibits an error floor in the high-SNR region, similar to the existing deep-learning-based coding schemes such as Deepcode and GBAF. We believe a key reason for this behavior is the difficulty of training the decoder in the high SNR region, where the error events are extremely rare. For instance, while training in the regime of BLER $10^{-9}$, most of the training batches (batch size $10^{5}$) are error-free and do not provide sufficient guidance for the gradient descent to learn useful information, leading to saturation in performance.
@@ VII-B Performance in moderate blocklength regime
We now present a modular way to scale LightCode to larger block lengths to enable a fair comparison against baselines with longer blocklengths. As a result of the curse of dimensionality, it is well known that the hardness of learning a code increases considerably with an increase in block length. As seen from results in Appendix Section A, directly increasing the blocklength for LightCode results in a poor BLER performance. GBAF [26] uses a computationally intensive transformer for performing block coding to support a length of $51$. To keep the complexity low, LightCode instead uses a short blocklength $K$ (e.g., $K=3$) and treats a block of $L$ bits as $l=\nicefrac{{L}}{{K}}$ independent sub-blocks. This modular approach of encoding $K$ bits at a time is motivated by ablation studies that demonstrate the negligible benefit of block coding in GBAF, presented in detail in Section VIII-A. Thus, the resulting BLER for a blocklength $L$ LightCode can be computed as
$p_{L}=1-(1-p_{K})^{l},$ (20)
where $p_{L}$ is the BLER for blocklength $L$ and $p_{K}$ is the BLER for blocklength $K$.
Finally, for a fair comparison against schemes with block length $K=50$, we refer to the results in Fig. 6 where we compute the BLER for blocklength $L=51$ for LightCode using the modular approach presented above. At a forward SNR of $-1.0$ dB, LightCode with blocklength $51$ has a BLER of $7\times 10^{-9}$ which is significantly smaller than Deepcode, DEFC, DRFC, and AttentionCode that use a block length of $50$, demonstrating the superior performance of LightCode even at moderately longer blocklengths. We note that the BLER $p_{L}$ approaches $1$ for very large values of $L$, but we restrict our study to the relatively short block length regime of $L<300$, where the performance is still superior to the baselines as demonstrated in Fig. 6. Further potential performance improvements at longer blocklengths may be achievable through a concatenated coding scheme that employs an outer block code, as explored in [6]; we leave this as a direction for future work.
Figure 6: By independently encoding sub-blocks of length $K=3$, LightCode provides a flexible way to encode for any blocklength $L$, providing a steady trade-off in BLER performance vs blocklength. Even at a block length of $50$, LightCode significantly outperforms all baseline methods shown in Fig. 5, which utilize a maximum block length of $51$.
@@ VII-C Coding for Channels with Noisy Feedback
Finally, we now consider the case of channels with noisy feedback i.e., $\sigma_{fb}^{2}>0$. While the assumption of noiseless feedback is easier to study, most practical feedback channels suffer from noise even when the receiver sending the feedback operates at high power. Because of this limitation, neither the SK nor GN schemes perform well. To address this, [6] introduces a linear feedback scheme that is implemented as an inner code to a concatenated code, which was found to be asymptotically optimal within the linear family of codes under AWGN forward channel [30]. More recently, [7] proposed using dynamic programming to improve the performance, which turns out to be a generalized version of SK. However, despite these improvements, the linearity of these schemes severely limits the performance that can be achieved.
Recalling the architecture of deep-learning-based schemes introduced in Section V, one of the interesting properties of these codes is their robustness to noise in the feedback channel. By taking advantage of the general architecture, injecting noise into the feedback channel during training is straightforward to make the encoder-decoder robust. Hence, learning-based schemes can help design a practically realizable class of codes that can be trained in a data-driven fashion and can provide gains in noiseless and noisy feedback settings.
Figure 7: Noisy feedback: Performance comparison against existing neural feedback codes for rate $\nicefrac{{3}}{{9}}$ and feedback SNR = $20$ dB. LightCode achieves BLER performance comparable to GBAF while utilizing only $\nicefrac{{1}}{{10}}^{\text{th}}$ the number of parameters.
@@ VIII Analysis
As evident from results in Section VII, symbol-by-symbol neural coding schemes such as LightCode can achieve performance comparable to block coding schemes such as GBAF. Consequently, we examine GBAF and LightCode in greater detail to systematically analyze their architecture, training process, and performance, with the aim of identifying the key factors contributing to their performance. Furthermore, we compare the memory and computational complexity of LightCode with existing schemes to quantify the gains. Finally, we provide an interpretation of the encoded representations of LightCode.
@@ VIII-A Ablation studies on GBAF
We investigate the contribution of the self-attention mechanism to the performance of GBAF by performing an ablation study to understand the compute vs performance trade-off better. First, the self-attention mechanism is disabled, and it is observed that this has no significant effect on the BLER performance. Next, both self-attention and positional encoding blocks are disabled, and the performance remains approximately the same. These results, provided in Table IV, demonstrate a surprising observation that the self-attention and PE modules, which are responsible for cross-symbol block coding, contribute only marginally to the performance of GBAF. This brings into question the value of block encoding and motivates us to find simpler designs that perform only symbol-by-symbol coding while providing competing performance.
@@
TABLE IV: Effect of block coding on performance of GBAF for rate $\nicefrac{{3}}{{9}}$ code with noiseless feedback. Positional encoding and self-attention modules have no noticeable effect on BLER performance.
@@ VIII-B Scaling laws: batch size
In section Section VII, we have demonstrated that it is possible to achieve performance comparable to GBAF with fewer parameters and lower compute complexity. We also hypothesized in Section VI-B that it is important to train the encoder using a very large batch size to accurately capture the statistics of the distribution so that the available power can be optimally allocated to the symbols with error at the receiver while reducing the power to the remaining symbols in a batch. In deep learning literature, it is well known that increasing the batch size can noticeably improve the performance of the neural network model [31]. To better understand the effect of batch size on LightCode, we perform a systematic study by training LightCode with different batch sizes. As noted in Fig. 8, at a batch size of $1.5\times 10^{3}$, LightCode has a BLER of $3.7\times 10^{-9}$, similar to that of GBAF (w/o BU). However, the BLER drops to $4\times 10^{-10}$, outperforming GBAF when the batch size increases to $5\times 10^{4}$ and beyond. We would like to note that for each batch size, hyperparameters such as learning rate have been optimized, and the model is trained until saturation to ensure the best possible performance.
@@
Using a lightweight network with a small number of parameters makes it feasible to train with a very large batch size, resulting in a significantly lower error rate.
@@ VIII-C Complexity and Throughput
We compare the total number of parameters in the encoder and decoder for a rate $\nicefrac{{3}}{{9}}$ LightCode against GBAF and Robust Coding within Table V. LightCode reduces the parameter count by more than 10×10$\times$10 × compared to GBAF and Robust Coding, resulting in notable savings in memory.
Scheme Total # parameters
@@
TABLE V: Rate $\nicefrac{{3}}{{9}}$ LightCode requires $<\nicefrac{{1}}{{10}}^{\text{th}}$ the number of parameters compared to GBAF and Robust Coding.
@@
TABLE VI: Throughput (symbols/sec) comparison. Rate $\nicefrac{{3}}{{9}}$ LightCode achieves up to 10×10$\times$10 × higher decoding throughput compared to GBAF and up to 171×171$\times$171 × higher decoding throughput in CPU mode compared to Robust Coding.
To compare the throughputs, we measure the inference time. We define the encoder throughput $T_{E}$ as the number of message blocks or symbols encoded per second and the decoder throughput $T_{D}$ as the number of message blocks or symbols decoded per second. Formally, for a rate $\nicefrac{{K}}{{D}}$ code, the throughput can be defined as
[…]
We measure the throughput in CPU mode on a AMD Ryzen Threadripper PRO 5975WX 32-Cores processor and in GPU mode using an NVIDIA GeForce RTX 4090, using a batch size of $10^{5}-10^{7}$. LightCode provides up to 10×10$\times$10 × higher encoding and decoding throughput than GBAF. Further, LightCode provides up to 171×171$\times$171 × higher decoding throughput in CPU mode and up to 10×10$\times$10 × higher throughput in GPU mode compared to Robust Coding, providing significant gains in latency, as shown in Table VI.
We note here that comparing throughputs for classical schemes against deep learning-based schemes is not straightforward and depends heavily on the implementation. Our implementation of LightCode uses PyTorch libraries, whereas PowerBlast and other classical schemes were implemented using NumPy libraries. Further, a significant computational bottleneck for the classical schemes is the necessity for demodulation after each round, the speed of which is heavily influenced by the implementation methodology. Hence, we restrict our comparison to the family of deep learning codes, where the latency primarily originates from the architecture complexity and the total number of parameters, and a fair comparison is feasible. However, it is clear that the total number of numerical operations in classical schemes is significantly lower than that of deep-learning-based schemes, and we acknowledge that the throughput of classical schemes can be significantly higher.
Complexity vs. BLER trade-off: While analytically quantifying the complexity of deep learning-based channel coding schemes with respect to target BLER would be very interesting, it is highly non-trivial and a widely open problem. Thus, we instead conduct an empirical study to analyze how the complexity of LightCode scales with target BLER. Empirically, we observed that reducing the complexity of the decoder while maintaining the same complexity for the encoder provided the best trade-off in performance vs complexity. In this experiment, we vary the target BLER and empirically find the required encoder-decoder dimensions for a rate $\nicefrac{{3}}{{9}}$ code at a forward SNR of $-1.0$ dB and a noiseless feedback channel. The results for the same are shown in Table VII, demonstrating an almost linear degradation of BLER in log scale with the number of parameters.
BLER Enc dimension Dec dimension Total # params
$4.5\times 10^{-10}$ 32 32 $7.3\times 10^{3}$
$3.4\times 10^{-9}$ 32 16 $4.7\times 10^{3}$
$5.2\times 10^{-9}$ 32 12 $4.3\times 10^{3}$
$2.8\times 10^{-8}$ 32 8 $3.9\times 10^{3}$
TABLE VII: Complexity vs BLER trade-off for rate $\nicefrac{{3}}{{9}}$ code at a forward SNR of $-1.0$ dB and noiseless feedback. BLER (in log scale) degrades almost linearly with a decrease in the number of parameters.